the average of All scores on 3 tests was 80. If the average of his scores on the first 2 tests was 77, what was the score on the third test?
Answer by hamsanash1981@gmail.com(151)

The average scores in 3 test is 80, then total score = 3* 80 => 240.
Similarly the total scores in 2 tests = 2*77 = 154
The third test score = 240- 154 = 86.